The Golden Ratio. The ancient Greeks thought that the most pleasing shape for a rectangle was one for which the ratio of the length to the width was 8 to 5, the golden ratio. If the length of the rectangular painting is 2 feet longer than its width, then for which dimensions would the length and the width have the golden ratio?
Answer by edjones(8007)

L/w=8/5 golden ratio
L=w+2
(w+2)/w=8/5 Proportion
5(w+2)=8w Multiply each side by the LCM (5w) to eliminate fractions.
5w+10=8w
3w=10
w=10'/3=3' 4"
L=5' 4"
.
Ed